# Integrated Vision and Language Lab, KAIST, South Korea

### v1

@@ 3.3 Finetuning: Multilingual Visual Speech Recognition
Even though we can directly perform multilingual VSR with the pre-trained model with discrete inputs, it is hard to outperform the model using continuous features. This is expected, as information is lost during the quantization process. However, as the pre-trained model has already learned how to model the multilingual visual pronunciation and generate languages, the finetuning with the continuous features is straightforward and more effective than direct training of the multilingual VSR model from scratch. For finetuning, we detach the unit embedding and linear layers, and attach the pre-trained mAV-HuBERT, as illustrated in Fig. 1(c). The model is trained end-to-end with the same objective function and data as the pre-training. In Sec. 4.3.5, we analyze the performances of multilingual VSR using the single proposed model and multiple previous monolingual models.
4 Experiment
4.1 Dataset
@@
Lip Reading Sentences 2 (LRS2) [47] is one of the largest English datasets for VSR. The dataset consists of 223 hours of training data collected from British TV shows.
@@
Lip Reading Sentences 3 (LRS3) [48] is a popular English VSR database. It has about 430 hours of video, and each video clip is collected from TED and TEDx. We evaluate the English VSR performances on LRS3.
@@
Multilingual TEDx (mTEDx) [72] is a multilingual dataset originally proposed for speech recognition and translation. The dataset provides 8 languages collected from TEDx talks. As the dataset also provides the video links of original talks, we download the video online. We remove the unavailable videos for VSR by referring to [18]. We use four languages, Spanish (Es), Italian (It), French (Fr), and Portuguese (Pt) for training and evaluating the developed VSR model, following [18, 19, 20].
@@
VoxCeleb2 [68] is a multilingual audio-visual dataset for speaker recognition [73]. This dataset has over 1 million utterances and contains 6,112 celebrities. For mAV-HuBERT training, we use the data corresponding to the following 9 languages, English (En), Spanish (Es), Italian (It), French (Fr), Portuguese (Pt), German (De), Russian (Ru), Arabic (Ar), and Greek (El), by identifying the language identity using Whisper [15]. For VSR training, as the dataset does not provide text annotations, we use the automatic labels of [39] for En, and [20] for Es, It, Fr, and Pt.
@@
Audio Visual Speech Dataset (AVSpeech) [69] is a large-scale audio-visual speech dataset. AVSpeech contains roughly 290k YouTube videos, and the total duration of these videos is 4700 hours. Since the dataset does not provide text annotations, we use the same strategy with the VoxCeleb2, using 9 languages for mAV-HuBERT and automatic labels for VSR.
@@
Table 2: Comparisons between English AV-HuBERT and the proposed multilingual AV-HuBERT in multilingual modeling. English (En) is validated on LRS3 and other languages are validated on mTEDx databases.
@@
4.2 Implementation Details
Preprocessing. The video is resampled to 25 fps. We detect the facial landmarks using RetinaFace [74], crop mouth regions using $96\times 96$ sizes of bounding box, and convert them into grayscale. For data augmentation, we randomly crop the video into $88\times 8$ 88 and horizontally flip it during training. The audio is resampled to 16kHz. In order to obtain the audio speech unit, we feed the audio into a multilingual trained HuBERT [31, 12] and cluster the extracted features with 1,000 token size. Finally, the audio speech unit is resampled to 25 fps to align with the sampling rate of the visual inputs. For the text, we construct one multilingual dictionary with 1,000 subword units by using SentencePiece tokenizer [75].
Architecture. Our mAV-HuBERT has the same architecture as the AV-HuBERT [37] large configuration. For visual speech unit-to-text translation, the model is composed of two unit embedding layers, one linear layer, one language embedding layer, 6 transformer encoders, and 6 transformer decoders. Each unit embedding layer embeds 1,000 tokens into 1,024 dimensions and the linear layer reduces the concatenated 2,048 dimensions into 1,024. The language token is embedded into a 1,024-dimensional feature by the language embedding layer. Each transformer layer has an embedding dimension of 1024, a feed-forward dimension of 4096, and 16 heads.
Table 3: Efficiency comparisons between previous VSR method and the proposed method. All numbers are measured using the same CPU and GPU (RTX 3090 24GB) environments. Test Acc means the subword-level prediction accuracy without beam search decoding.
@@ 4.2 Implementation Details
Fig. 2: Finetuning efficiency comparison between the proposed pre-training scheme and without the pre-training.
Fig. 3: Visualization of speech units. Each boundary represents a single unit and the same color represents the same phoneme or phoneme family. (a) Audio speech unit. (b) Visual speech unit.
Table 4: Speaker verification results (EER) comparisons using different input representations.
@@
Training. For training mAV-HuBERT, we follow the original AV-HuBERT [37] and train it with the masked prediction task. For the prediction target, we use 1,000 clusters extracted from multilingual trained HuBERT [31, 12]. We train the model for 350k steps using 64 3090 RTX GPUs. For pre-training the proposed model (i.e., visual speech unit-to-text translation), we train the model for 11 epochs with a tri-stage learning rate scheduler and 32 GPUs. For the progressive masking $\mathcal{M(\cdot)}$, we set $p$ as 0 for the first 10% of training. From 10% to 70% of training, we linearly increase the masking ratio $p$ from 0 to 100. After 70% of training, $p$ is set to 100 so that only visual speech units are used. We finetune the model with the continuous features for 8 epochs using 32 GPUs. For all experiments, the Adam optimizer [76] is used. For beam search decoding, we use a beam width chosen from {20, 25, 30, 35} and a length penalty of 0. The detailed training configuration can be found in the supplementary.
4.3 Experimental Results
4.3.1 Effectiveness of mAV-HuBERT in Multilingual Modeling
In order to verify the effectiveness of mAV-HuBERT in modeling multilingual speech, we compare our mAV-HuBERT with the English-trained AV-HuBERT. To this end, we finetune the pre-trained mAV-HuBERT and AV-HuBERT on 5 languages (i.e., En, Es, It, Fr, Pt) using LRS3 and mTEDx databases. The multilingual VSR performances of the two models are shown in Table 2. The results show that when performing multilingual VSR with a single model, the mAV-HuBERT is much more effective than the English-only trained AV-HuBERT. Specifically, for Es, It, Fr, and Pt, the proposed mAV-HuBERT outperforms the AV-HuBERT with large margins of over 10% WERs. English VSR achieves better performance when utilizing English-trained AV-HuBERT. This is related to the previous observations in [77, 78], where an increasing number of languages can enhance cross-lingual performance up to a certain threshold, beyond which performance drops may occur. As the proposed mAV-HuBERT shows better performances in modeling multilingual visual speech, we extract the visual speech unit by using the mAV-HuBERT.
@@ 4.3.1 Effectiveness of mAV-HuBERT in Multilingual Modeling
Table 5: The effectiveness of each proposed component. By substituting each component from the full model, we evaluate their effectiveness in multilingual VSR.
@@
4.3.2 Efficiency of Visual Speech Units and Raw Inputs
We compare the batch size, training iteration time, and total training time between the standard VSR method that uses raw video as inputs and the proposed method. Table 3 shows the comparison results. By using the visual speech units as inputs in pre-training, we can increase the batch size sixfold and reduce the training iteration time by approximately half. Consequently, we can expedite training by a factor of about 12 compared to previous VSR training methods. The total training time for pre-training the model for 11 epochs amounts to 6.6 hours, whereas the conventional VSR model requires 52.5 hours for 8 epochs of training. Furthermore, when fine-tuning the pre-trained model, we can achieve better performance compared to the standard VSR model, with just 5 epochs of fine-tuning taking 34.9 hours. As a result, we can greatly boost training time even if considering both the pre-training and finetuning stages, compared to the standard VSR training. To see this more intuitive, we examine the subword-level accuracy changes during fine-tuning, comparing the proposed method with the model without pre-training. Fig. 2 shows the learning curve of the two models. After the proposed pre-training using visual speech units, the finetuning on the continuous features is effective so that we can achieve better performance even with much fewer epochs than the model without pre-training. It’s worth noting that we can significantly reduce the pre-training time with the proposed method; 1 epoch requires just 0.6 hours in pre-training, whereas the standard VSR training demands 6.6 hours for 1 epoch.
Table 6: Multilingual VSR performance comparisons. As there is no prior work that can perform multilingual VSR with a single model, we train AV-HuBERT to perform multilingual VSR.
@@ 4.3.2 Efficiency of Visual Speech Units and Raw Inputs
Fig. 4: Visualization of video frames corresponding to visual speech units. Each number indicates an index of visual speech unit.
4.3.3 Analyzing the Characteristics of Visual Speech Units
To understand which information is held by the visual speech units, we analyze them using 1) phoneme mapping visualization [32] and 2) speaker recognition. Firstly, following [32], we visualize the phonetic information of the audio speech unit and visual speech unit which are obtained from mAV-HuBERT. Therefore, only audio inputs are used to extract audio speech units and video inputs for visual speech units. We set the cluster centroid as each row of the weight of the final layer of mAV-HuBERT (i.e., classifier for 1,000 units), so that the same boundary can be obtained for different modalities. Fig. 3 displays the visualization results of the 200 units out of 1,000 that appeared most frequently. We can confirm that each unit in both the audio speech units and visual speech units contains distinctive linguistic information (i.e., phoneme or viseme). By comparing the audio speech unit and visual speech unit, we can find that the homophenes which refer to the different pronunciations having the same lip movement (e.g., name, tame, dame) are confusingly represented in visual speech units [56]. For example, some units representing ‘n’ in audio speech units are changed to ‘d’ or ‘t’ in visual speech units. Moreover, we can find that compared to audio speech units, more visual speech units are pointing to the vowel (i.e., blue-colored area), which shows the ambiguity of lip movements compared to the audio modality. These are the natural results reflecting the characteristics of different speech modalities, and should not be taken as an indication of the inadequacy of the visual modality in speech modeling. In addition, we visualize some video frames corresponding to each visual speech unit in Fig. 4, illustrating that similar lip movements are consistently mapped to the same index regardless of pose variations and speakers. For example, the 648-th visual speech unit represents lip frames related to the viseme ‘a’, and the 912-th visual speech unit corresponds to the viseme ‘o’. Through the visualization, we can confirm that our visual speech units contain the linguistic information, the viseme, which enables us to pre-train the model to build the knowledge of visual speech modeling.
Table 7: VSR performance comparisons with the state-of-the-art monolingual VSR methods. Please note that the proposed method utilizes a single multilingual model while the other methods utilize language-specific models. Best and second-best scores are bolded and underlined.
@@ 4.3.3 Analyzing the Characteristics of Visual Speech Units
En Ma et al[18] - 1,459 31.5
Prajwal et al[53] - 2,676 30.7
AV-HuBERT [37] 1,759 433 28.6
VATLM [79] 1,759 433 28.4
Haliassos et al[38] 1,759 433 27.8
AKVSR [80] 1,759 433 27.6
Auto-AVSR [39] - 3,448 20.5
@@
Secondly, we analyze how much degree the speech units contain non-linguistic information through speaker recognition. To this end, we use raw audio speech, audio/visual speech features extracted from mAV-HuBERT, and audio/visual speech units as inputs to perform the speaker recognition. For the model, we follow [81] and train the model on each input. The speaker verification results (Equal Error Rate; EER) are shown in Table 4. The random prediction yields 50% EER and 0% EER means a perfect system. When we use raw audio as input, the system can almost perfectly distinguish the input speakers with 2.38% EER. When we use the features of the pre-trained mAV-HuBERT, the verification performance is dropped to 14.96% and 19.42% on audio and visual modalities, respectively. This shows that the masked prediction of AV-HuBERT [37] forces the model to learn linguistic information while somewhat discarding the speaker information. Finally, when we discretize the speech features of mAV-HuBERT and obtain the speech units, we can greatly suppress the speaker information. The performance is dropped to 28.84% by using the audio speech unit and 32.74% by using the visual speech unit. Through the above two experiments, we can confirm that the quantization process averages out the speaker effects (i.e., non-linguistic information) in the speech representations while maintaining the content. This enables us to build a speech representation model by using speech units as input.
4.3.4 Effectiveness of Each Proposed Component
To confirm the effectiveness of each proposed component, we substitute each component from the proposed method. The ablation results are shown in Table 5. The performance of ‘--Unit Pretraining’ is obtained by directly finetuning the proposed mAV-HuBERT on the multilingual video-text without using visual speech units. In this case, the overall VSR performances are dropped especially for non-English languages. Moreover, we require more training times compared to the proposed method as shown in Table 3. When we do not utilize curriculum learning (i.e., ‘--Curriculum’), the performance dramatically decreases and even shows worse performances than the without pre-training method in some languages. This shows that directly performing the visual speech unit to text translation from scratch is challenging to the model in finding the optimal points. Therefore, the proposed curriculum learning is crucial when learning from the visual speech units. These results also coincide with the previous methods utilizing multi-modal complementary in VSR training [55, 37, 34, 35, 56, 80]. Finally, when we directly perform multilingual VSR with the pre-trained model (i.e., with visual speech units), we cannot achieve satisfactory performances. Therefore, finetuning with the continuous features for a few epochs should be performed to maximize the pre-trained knowledge.
4.3.5 Multilingual Visual Speech Recognition with a Single Model
We validate the effectiveness of the proposed multilingual VSR method by comparing it with 1) the multilingual VSR model and 2) the monolingual VSR model. Since there is no prior work exploring multilingual VSR with a single model, we train AV-HuBERT [37] to perform multilingual VSR and set it as the baseline. Moreover, since the previous non-English VSR methods are language-specific, we compare the performance of our model with multiple monolingual models.
@@ 4.3.5 Multilingual Visual Speech Recognition with a Single Model
Comparison with multilingual VSR method. Table 6 shows the performance comparison results of multilingual VSR methods. Both the AV-HuBERT and the proposed method are finetuned on 4,545 hours of multilingual video-text paired data. The proposed method outperforms the AV-HuBERT for all languages except English. In particular, the proposed method demonstrates significantly improved performance for Es, It, Fr, and Pt, with gains of more than 4% WERs. For the high-resource language En, the proposed method achieves similar performance with AV-HuBERT but slightly falls behind, as we have seen in Sec. 4.3.1. Considering multilingualism, the results confirm that the proposed strategy using visual speech units is much more effective in building multilingual VSR models by achieving new state-of-the-art multilingual VSR performances.
@@
Comparison with monolingual VSR method. Here, we compare the proposed multilingual VSR performances with the previous state-of-the-art monolingual VSR methods. Please note that the proposed method utilizes a single trained model across the languages, while different methods utilize multiple language-specific VSR models. The results are shown in Table 7. By comparing the recent state-of-the-art method [20], we outperform it in 3 languages, except Portuguese (Pt). In the English VSR, our method achieves 24.4% WER while the current previous method [39] achieves 20.5% WER. Therefore, the proposed multilingual VSR model achieves the best score in Es, It, and Fr, and the second-best score in En and Pt with a single-trained model. As discussed in [15], multilingual models benefit more from scale than monolingual models. Thus, scaling up the model size may improve the multilingual VSR performances overall. Through the comparisons, we can confirm the effectiveness of the proposed method in multilingual VSR by outperforming and achieving comparable results with the previous language-specific VSR methods.
5 Conclusion

### v2

@@ 3.3. Finetuning: Multilingual Visual Speech Recognition
Even though we can directly perform multilingual VSR with the pre-trained model with discrete inputs, it is hard to outperform the model using continuous features. This is expected, as information is lost during the quantization process. However, as the pre-trained model has already learned how to model the multilingual visual pronunciation and generate languages, the finetuning with the continuous features is straightforward and more effective than direct training of the multilingual VSR model from scratch. For finetuning, we detach the unit embedding and linear layers, and attach the pre-trained mAV-HuBERT, as illustrated in Fig. 1(c). The model is trained end-to-end with the same objective function and training data as the pre-training. In Sec. 4.3.5, we analyze the performances of multilingual VSR using the single proposed model and multiple previous monolingual models.
4. Experiment
4.1. Dataset
@@
Lip Reading Sentences 2 (LRS2) (Chung et al., 2017) is one of the largest English datasets for VSR. The dataset consists of 223 hours of training data collected from British TV shows.
@@
Lip Reading Sentences 3 (LRS3) (Afouras et al., 2018a) is a popular English VSR database. It has about 430 hours of video, and each video clip is collected from TED and TEDx. We evaluate the English VSR performances on LRS3.
@@
Multilingual TEDx (mTEDx) (Elizabeth et al., 2021) is a multilingual dataset originally proposed for speech recognition and translation. The dataset provides 8 languages collected from TEDx talks. As the dataset also provides the video links of original talks, we download the video online. We remove the unavailable videos for VSR by referring to (Ma et al., 2022a). We use four languages, Spanish (Es), Italian (It), French (Fr), and Portuguese (Pt) for training and evaluating the developed VSR model, following (Ma et al., 2022a; Kim et al., 2023c; Yeo et al., 2023b).
@@
VoxCeleb2 (Chung et al., 2018) is a multilingual audio-visual dataset for speaker recognition (Jung et al., 2022). This dataset has over 1 million utterances and contains 6,112 celebrities. For mAV-HuBERT training, we use the data corresponding to the following 9 languages, English (En), Spanish (Es), Italian (It), French (Fr), Portuguese (Pt), German (De), Russian (Ru), Arabic (Ar), and Greek (El), by identifying the language identity using Whisper (Radford et al., 2023). For VSR training, as the dataset does not provide text annotations, we use the automatic labels of (Ma et al., 2023) for En, and (Yeo et al., 2023b) for Es, It, Fr, and Pt.
@@
Audio Visual Speech Dataset (AVSpeech) (Ephrat et al., 2018) is a large-scale audio-visual speech dataset. AVSpeech contains roughly 290k YouTube videos, and the total duration of these videos is 4700 hours. Since the dataset does not provide text annotations, we use the same strategy with the VoxCeleb2, using 9 languages for mAV-HuBERT and automatic labels for VSR.
4.2. Implementation Details
Preprocessing. The video is resampled to 25 fps. We detect the facial landmarks using RetinaFace (Deng et al., 2020), crop mouth regions using $96\times 96$ sizes of bounding box, and convert them into grayscale. For data augmentation, we randomly crop the video into $88\times 8$ 88 and horizontally flip it during training. The audio is resampled to 16kHz. In order to obtain the audio speech unit, we feed the audio into a multilingual trained HuBERT (Hsu et al., 2021; Lee et al., 2022) and cluster the extracted features with 1,000 token size. Finally, the audio speech unit is resampled to 25 fps to align with the sampling rate of the visual inputs. For the text, we construct one multilingual dictionary with 1,000 subword units by using SentencePiece tokenizer (Kudo and Richardson, 2018).
Architecture. The mAV-HuBERT has the same architecture as the AV-HuBERT (Shi et al., 2021) large configuration and only differs in pre-training data. For visual speech unit-to-text translation, the model is composed of two unit embedding layers, one linear layer, one language embedding layer, 6 transformer encoders, and 6 transformer decoders. Each unit embedding layer embeds 1,000 tokens into 1,024 dimensions and the linear layer reduces the concatenated 2,048 dimensions into 1,024. The language token is embedded into a 1,024-dimensional feature by the language embedding layer. Each transformer layer has an embedding dimension of 1024, a feed-forward dimension of 4096, and 16 heads.
@@ 4.2. Implementation Details
Table 2. Comparisons between English AV-HuBERT and the multilingual AV-HuBERT in multilingual VSR. English (En) is validated on LRS3 and other languages (Es, It, Fr, Pt) are validated on mTEDx databases.
@@
Table 3. Efficiency comparisons between previous VSR method and the proposed method. All numbers are measured using the same CPU and GPU (RTX 3090 24GB) environments. Test Acc means the subword-level prediction accuracy without beam search decoding.
@@
Training. For training mAV-HuBERT, we follow the original AV-HuBERT (Shi et al., 2021) and train it with the masked prediction task. For the prediction target, we use 1,000 clusters extracted from multilingual trained HuBERT (Hsu et al., 2021; Lee et al., 2022). We train the model for 350k steps using 64 3090 RTX GPUs. For pre-training the proposed model (i.e., visual speech unit-to-text translation), we train the model for 11 epochs with a tri-stage learning rate scheduler and 32 GPUs. For the progressive masking $\mathcal{M(\cdot)}$, we set $p$ as 0 for the first 10% of training. From 10% to 70% of training, we linearly increase the masking ratio $p$ from 0 to 100. After 70% of training, $p$ is set to 100 so that only visual speech units are used. We finetune the model with the continuous features for 8 epochs using 32 GPUs. For all experiments, the Adam optimizer (Kingma and Ba, 2015) is used. For beam search decoding, we use a beam width chosen from {20, 25, 30, 35} and a length penalty of 0. The detailed training configuration can be found in the supplementary.
4.3. Experimental Results
4.3.1. Effectiveness of mAV-HuBERT in Modeling Multilingual Speech
Before extracting the visual speech units by using SSL visual speech models, we need to confirm which model is best suitable for our multilingual VSR purposes. Therefore, we compare the performances of English-trained AV-HuBERT and multilingual-trained AV-HuBERT (mAV-HuBERT). To this end, we finetune the pre-trained mAV-HuBERT and AV-HuBERT on 5 languages (i.e., En, Es, It, Fr, Pt) by merging LRS3 and mTEDx databases. The multilingual VSR performances of the two models are shown in Table 2. The results show that when performing multilingual VSR with a single model, the mAV-HuBERT is much more effective than the English-only trained AV-HuBERT. Specifically, for Es, It, Fr, and Pt, the mAV-HuBERT outperforms the AV-HuBERT with large margins of over 10% WERs. For English, English-only trained AV-HuBERT achieves better performance, and this tendency is also observed in Radford et al. (2023); Increasing language diversity without scaling the model size, the performances for major languages can be lower than the monolingual model. As the mAV-HuBERT shows better performances in modeling multilingual visual speech, we extract the visual speech unit by using the mAV-HuBERT.
Figure 2. Finetuning efficiency comparison between the proposed pre-training scheme and without the pre-training.
Figure 3. Visualization of speech units. Each boundary represents a single unit and the same color represents the same phoneme or phoneme family. (a) Audio speech unit. (b) Visual speech unit.
Table 4. Speaker verification results (EER) comparisons using different input representations.
@@ 4.3.1. Effectiveness of mAV-HuBERT in Modeling Multilingual Speech
4.3.2. Efficiency Comparison between Visual Speech Units and Raw Inputs
To confirm the efficiency of using visual speech units instead of raw videos as inputs, we compare the batch size, training iteration time, and total training time between the standard VSR method that uses raw video as inputs and the proposed method. Table 3 shows the comparison results. By using the visual speech units as inputs in pre-training, we can increase the batch size sixfold and reduce the training iteration time by approximately half. Consequently, we can expedite training by a factor of about 12 compared to previous VSR training methods. The total training time for pre-training the model for 11 epochs amounts to 6.6 hours, whereas the conventional VSR model requires 52.5 hours for 8 epochs of training. Furthermore, when fine-tuning the pre-trained model, we can achieve better performance compared to the standard VSR model, with just 5 epochs of fine-tuning taking 34.9 hours. As a result, we can greatly boost training time even if considering both the pre-training and finetuning stages, compared to the standard VSR training. To see this more intuitive, we examine the subword-level accuracy changes during fine-tuning, comparing the proposed method with the model without pre-training. Fig. 2 shows the learning curve of the two models. After the proposed pre-training, the finetuning on the continuous features is effective so that we can achieve better performance even with much fewer epochs than the model without pre-training. It’s worth noting that we can significantly reduce the pre-training time with the proposed method; 1 epoch requires just 0.6 hours in pre-training, whereas the standard VSR training demands 6.6 hours for 1 epoch.
4.3.3. Analyzing Visual Speech Units
To understand which information is held by the visual speech units, we analyze them using 1) phoneme mapping visualization (Sicherman and Adi, 2023) and 2) speaker recognition. Firstly, following (Sicherman and Adi, 2023), we visualize the phonetic information of the audio speech unit and visual speech unit which are obtained from mAV-HuBERT. Therefore, only audio inputs are used to extract audio speech units and video inputs for visual speech units. We set the cluster centroid as each row of the weight of the final layer of mAV-HuBERT (i.e., classifier for 1,000 units), so that the same boundary can be obtained for different modalities. Fig. 3 displays the visualization results of the 200 units out of 1,000 that appeared most frequently. We can confirm that each unit in both the audio speech units and visual speech units contains distinctive linguistic information (i.e., phoneme or viseme). By comparing the audio speech unit and visual speech unit, we can find that the homophenes which refer to the different pronunciations having the same lip movement (e.g., name, tame, dame) are confusingly represented in visual speech units (Kim et al., 2022). For example, some units representing ‘n’ in audio speech units are changed to ‘d’ or ‘t’ in visual speech units. Moreover, we can find that compared to audio speech units, more visual speech units are pointing to the vowel (i.e., blue-colored area), which shows the ambiguity of lip movements compared to the audio modality. These are the natural results reflecting the characteristics of different speech modalities, and should not be taken as an indication of the inadequacy of the visual modality in speech modeling. In addition, we visualize some video frames corresponding to each visual speech unit in Fig. 4, illustrating that similar lip movements are consistently mapped to the same index regardless of pose variations and speakers. For example, the 648-th visual speech unit represents lip frames related to the viseme ‘a’, and the 912-th visual speech unit corresponds to the viseme ‘o’. Through the visualization, we can confirm that our visual speech units contain the linguistic information, the viseme, which enables us to pre-train the model to build the knowledge of visual speech modeling.
Secondly, we analyze how much degree the speech units contain non-linguistic information through speaker recognition. To this end, we use raw audio speech, audio/visual speech features extracted from mAV-HuBERT, and audio/visual speech units as inputs to perform the speaker recognition. For the model, we utilize a pre-trained speaker recognition model of (Desplanques et al., 2020) and train the model with different inputs. To match the input size with the model, we use one additional embedding layer for both audio speech units and visual speech units. When we use the speech features from mAV-HuBERT, we utilize one additional linear layer for both audio speech features and visual speech features. For training, we utilize data less than 20 seconds in VoxCeleb2 (Chung et al., 2018) dev set. The speaker verification is performed on the test data of VoxCeleb2 having less than 20 seconds, where each test sample is assigned one positive sample and one negative sample. Therefore, 27,816 positive pairs and negative pairs are utilized respectively, thus random prediction yields 50% EER.
Figure 4. Example frames corresponding to each visual speech unit. The numbers indicate the classes of visual speech units.
The speaker verification results (Equal Error Rate; EER) are shown in Table 4. When we use raw audio as input, the system can almost perfectly distinguish the input speakers with 2.38% EER. When we use the features of the pre-trained mAV-HuBERT, the verification performance is dropped to 14.96% and 19.42% on audio and visual modalities, respectively. This shows that the masked prediction of AV-HuBERT (Shi et al., 2021) forces the model to learn linguistic information while somewhat discarding the speaker information. Finally, when we discretize the speech features of mAV-HuBERT and obtain the speech units, we can greatly suppress the speaker information. The performance is dropped to 28.84% by using the audio speech unit and 32.74% by using the visual speech unit. Through the above two experiments, we can confirm that the quantization process averages out the speaker effects (i.e., non-linguistic information) in the speech representations while maintaining the content. This enables us to build a speech representation model by using speech units as input.
@@ 4.3.3. Analyzing Visual Speech Units
Table 5. The effectiveness of each proposed component (WER, %). By substituting each component from the full model, we evaluate their effectiveness in multilingual VSR.
@@
4.3.4. Effectiveness of Each Proposed Component
To confirm the effectiveness of each proposed component, we substitute each component from the proposed method. The ablation results are shown in Table 5. The performance of ‘--Unit Pretraining’ is obtained by directly finetuning the mAV-HuBERT on the multilingual video-text without using visual speech units. In this case, the overall VSR performances are dropped especially for non-English languages. Moreover, we require more training times compared to the proposed method as shown in Table 3. When we do not utilize curriculum learning (i.e., ‘--Curriculum’), the performance dramatically decreases and even shows worse performances than the without pre-training method in some languages. This shows that directly performing the visual speech unit to text translation from scratch is challenging to the model in finding the optimal points. Therefore, the proposed curriculum learning is crucial when learning from the visual speech units. These results also coincide with the previous methods utilizing multi-modal complementary in VSR training (Afouras et al., 2020; Shi et al., 2021; Zhao et al., 2020; Ren et al., 2021; Kim et al., 2022; Yeo et al., 2024). Finally, when we directly perform multilingual VSR with the pre-trained model (i.e., with visual speech units), overall performances are decreased as fine information is lost during the quantization process. Therefore, finetuning with the continuous features for a few epochs should be performed to maximize the pre-trained knowledge.
Table 6. Multilingual VSR performance (WER, %) comparisons. As there is no prior work that can perform multilingual VSR with a single model, we train AV-HuBERT to perform multilingual VSR.
@@ 4.3.4. Effectiveness of Each Proposed Component
4.3.5. Multilingual Visual Speech Recognition with a Single Trained Model
We validate the effectiveness of the proposed multilingual VSR method by comparing it with 1) the multilingual VSR model and 2) the monolingual VSR model. Since there is no prior work exploring multilingual VSR with a single model, we train AV-HuBERT (Shi et al., 2021) to perform multilingual VSR and set it as our baseline. Moreover, since the previous non-English VSR methods are language-specific, we compare the performance of our model with multiple monolingual models.
Table 7. VSR performance comparisons with the previous VSR methods. Please note that the proposed method utilizes a single model while the other methods utilize multiple models. Best and second-best scores are bolded and underlined.
@@ 4.3.5. Multilingual Visual Speech Recognition with a Single Trained Model
En Ma et al. (2022a) - 1,459 31.5
Prajwal et al. (2022) - 2,676 30.7
AV-HuBERT (Shi et al., 2021) 1,759 433 28.6
VATLM (Zhu et al., 2023) 1,759 433 28.4
Haliassos et al. (2022) 1,759 433 27.8
AKVSR (Yeo et al., 2024) 1,759 433 27.6
Auto-AVSR (Ma et al., 2023) - 3,448 20.5
@@
Comparison with multilingual VSR method. Table 6 shows the performance comparison results of multilingual VSR methods. Both the AV-HuBERT and the proposed method are finetuned on 4,545 hours of multilingual video-text paired data. The proposed method outperforms the AV-HuBERT for all languages except English. In particular, the proposed method demonstrates significantly improved performance for Es, It, Fr, and Pt, with gains of more than 4% WERs. For the high-resource language En, the proposed method achieves similar performance with AV-HuBERT but slightly falls behind. Considering multilingualism, the results confirm that the proposed VSR framework using visual speech units is much more effective in building multilingual VSR models by achieving new state-of-the-art performances. It is also worth noting that the proposed method is more efficiently trainable with the proposed pre-training strategy as discussed in Sec. 4.3.2.
@@
Comparison with monolingual VSR method. Here, we compare the proposed multilingual VSR performances with the previous state-of-the-art monolingual VSR methods. Please note that the proposed method utilizes a single trained model across the languages, while different methods utilize multiple language-specific VSR models. The results are shown in Table 7. By comparing with the recent state-of-the-art method (Yeo et al., 2023b), we outperform it in 3 languages Spanish (Es), Italian (It), and French (Fr), by 3.0%, 2.1%, and 3.1% WER, respectively. In the English VSR, our method achieves 24.4% WER while the current previous method (Ma et al., 2023) achieves 20.5% WER. As discussed in (Radford et al., 2023), the performance of high-resource languages can be further improved by scaling the model size to accommodate the diversity of languages. As a result, the proposed multilingual VSR model achieves the best score in Es, It, and Fr, and the second-best score in En and Pt with a single-trained model. Through the comparisons, we can confirm not only the computation efficiency but also the effectiveness of the proposed method in multilingual VSR by outperforming and achieving comparable results with the previous language-specific VSR methods.
5. Conclusion